Problem 3 Find v.lt) a) Using convolution integral b) Using transfer function c) when us = u(t) +

USE LAPLACE TRANSFORM
Answer #1

For Volt • 1 VyC+3) = {tucu) - [uct23] . Tau Laplace Frand fозы ең чы yoxsa za 20 circait m Vco + + . Grounded the lower refeCurrenti Pa wao P = VA a UA= is - • current through capacirog. say ices Pic0 V4 +61 las voltage acron capacitoy, Voco voce =• Peso = 325²49879 This Peso pur in equation ® Volso = (28f6Jx Us : 32547985+41 Diese i voro (831 * V 165²449572 Now put u -op. 168t 428.42 = (s+0.04)+3.02) Now Volso = (5+3/2 --287 لاک shesf004)(578.00 L Use partial fraction Voc&o : _ A f + S sto.0